OpenStudy (anonymous):

fifthroot(256x^17y^10). Simplify by factoring assume that all expressions under the radical represent nonnegative numbers. (I think I may have typed the expression wrong, if anyone knows, please let me know.)

OpenStudy (anonymous):

fifthroot(256x^17y^10)= fifthroot(32*8*x^15*x^2*(y^2)^5)= fifthroot(2^5*8*(x^3)^5*x^2*(y^2)^5)=2x^3y^2fifthroot(8x^2)=\[2x ^{3}y ^{2}\sqrt[5]{8x ^{2}}\].